[Brahms-l] Shift report 20040320 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Sat Mar 20 2004 - 23:51:03 EST
------------------------------------------------------------
Date: Sat Mar 20 23:51:01 2004, Shift: 16:00-24:00
Supervisor: EBJ, with: BN
Log book#15 pages 92 - 92

Shift Summary:
------------------------------------------------------------
1. We has a lot of good beam for the entire shift.

2. We completed tge data set for B1219 FS (1M trig6), and we are working on A1219 FS now.

3. At the beginning of the shift the DAQ was not responding well, and we have to restart the ioserver a number of times before the DAQ began collecting data.
  We noticed that gas flow rates for the TPC were from 0.2 to 0.4 higher than the default values.
  Also TOFW channel 110 seems to have problems.  I do not see it in the distribution.  See Elog.  
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Mar 20 23:51:20 2004

This archive was generated by hypermail 2.1.8 : Sat Mar 20 2004 - 23:51:43 EST